OpenStudy (egracer):

Sandi tracks her calories burned during water aerobics class. The number of calories she burns is expressed by the function c(t) = 350t, where t is the number of hours spent doing water aerobics. To burn more calories, Sandi wears ankle weights during her class. The number of calories she burns while wearing ankle weights is expressed by the function b(c) = 1.2c, where c is the number of calories burned doing water aerobics without weights. Which of the following composite functions expresses the calories, as a function of time, Sandi burns while doing water aerobics with ankle weights?

OpenStudy (egracer):

the options are b[c(t)] = 351.2t b[c(t)] = 420t c[b(t)] = 351.2t c[b(t)] = 420t

OpenStudy (egracer):

Open study kind of sucks nobody ever helps me..

OpenStudy (egracer):

Is it C????? thats what i got through by calculations

OpenStudy (egracer):

@ganeshie8 @Hero @phi

Vocaloid (vocaloid):

first find b(t) if b(c) = 1.2c, then what is b(t)?

OpenStudy (egracer):

b(t)=351.2 @Vocaloid

Vocaloid (vocaloid):

uh, not quite all we're doing is replacing "c" with "t" so, if b(c) is 1.2c, then what is b(t)?

Vocaloid (vocaloid):

hint: all you need to do is replace "c" with "t"

OpenStudy (egracer):

1.2t??

Vocaloid (vocaloid):

right now, we have b(t) = 1.2t now we need to find c(b(t)) if c(t) = 350t, then what is c(b(t))? hint: replace "t" with "1.2t"

OpenStudy (egracer):

350c(1.2t)

OpenStudy (egracer):

im not really sure I'm a little confused

Vocaloid (vocaloid):

c(t) = 350t c(b(t)) = c(1.2t) = ?

OpenStudy (egracer):

350(1.2)?

Vocaloid (vocaloid):

almost... if c(t) is 350 times t, then c(1.2t) is 350 times ____ fill in the blank

OpenStudy (egracer):

that is so confusing could you break it down some more its really confusign me

Vocaloid (vocaloid):

c(t) = 350*t c(1.2t) = 350*___ fill in the blank

OpenStudy (egracer):

350 times 1.2 thats all i can think of

Vocaloid (vocaloid):

almost, you're missing something important

OpenStudy (egracer):

the t?

Vocaloid (vocaloid):

yes!

OpenStudy (egracer):

so its 350*1.2t

Vocaloid (vocaloid):

yeah, and what do you get when you multiply?

OpenStudy (egracer):

420t

OpenStudy (egracer):

so would it be D?

Vocaloid (vocaloid):

ah, it should be B, my mistake b(c(t)) = 420t

OpenStudy (egracer):

thank you so much

OpenStudy (egracer):

EVERYONE B WAS CORRECT.